OpenStudy (anonymous):

I am stuck on a complex rational problem solve, simplify result. leave in factored form x-2/x+2 + x-1/x+1 / x/x+1 - 2x-3/x

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

\[\frac{x-2}{x+2} + \frac{x-1}{x+1 }\div (\frac{x}{x+1} - \frac{2x-3}{x})\]

OpenStudy (anonymous):

yeah that is the problem, thank you

OpenStudy (anonymous):

is that it? hard to read

OpenStudy (anonymous):

holy crap. your math teacher must really hate you.

OpenStudy (anonymous):

i was trying to type it the way you do but i don't i can't find a way to write it with the equation tool

OpenStudy (anonymous):

i do not see a snap way to do this, so you are going to have to add the first two terms, subtract the last two and then invert and multiply. i guess we just have to grind it out

OpenStudy (anonymous):

ready?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok first \[\frac{x-2}{x+2}+\frac{x-1}{x+1}=\frac{(x-2)(x+1)+(x+2)(x-1)}{(x+2)(x+1)}\]

OpenStudy (anonymous):

lets leave denominator in factored form. numerator is \[x^2-x-2+x^2+x-2=2x^2\] so we have for the first part \[\frac{2x^2}{(x+2)(x+1)}\]

OpenStudy (anonymous):

ok i follow

OpenStudy (anonymous):

now \[\frac{x}{x+1}-\frac{2x-3}{x}\] \[=\frac{x^2-(2x-3)(x+1)}{x(x+1)}\]

OpenStudy (anonymous):

again we leave denominator in factored form and multiply out in the top \[x^2-(2x^2-x-3)=x^2-2x^2+x+3=-x^2+x+3\]

OpenStudy (anonymous):

you should check my algebra but i think it is right. so second part is \[\frac{-x^2+x+3}{x(x+1)}\]

OpenStudy (anonymous):

i am trying to see where the \[x^{2}\] came from

OpenStudy (anonymous):

\[\frac{a}{b}-\frac{c}{d}=\frac{ac-bd}{bd}\] and here both a and c are x

OpenStudy (anonymous):

i mean both a and d are x

OpenStudy (anonymous):

still not done. now we have to invert and multiply

OpenStudy (anonymous):

ok i see

OpenStudy (anonymous):

\[\frac{2x^2}{(x+2)(x+1)}\times \frac{x(x+1)}{-x^2+x+3}\]

OpenStudy (anonymous):

you can cancel a factor of x+1 from top and bottom that is about it. unless i made some other mistake.

OpenStudy (anonymous):

i get \[\frac{2x^3}{(x+2)(-x^2+x+3)}\]

OpenStudy (anonymous):

check my algebra, but i think it is right. the method is certainly right

OpenStudy (anonymous):

the back of my book , shows -2x^2(x^2-2) for the top part. maybe my book is wrong

OpenStudy (anonymous):

it is entirely possible that i made a mistake somewhere.

OpenStudy (anonymous):

there is a mistake in the first fraction. it should come out to \[2x ^{2}-4\]-------- (x+2)(x+1)

OpenStudy (anonymous):

how do you like that ! you are right. it is \[x^2-4\] for the first one